# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

### v4

@@ Abstract
We propose Temporal Conformal Prediction (TCP), a distribution‐free framework for constructing well‐calibrated prediction intervals in nonstationary time series. TCP couples a modern quantile forecaster with a split‐conformal calibration layer on a rolling window and, in its TCP–RM variant, augments the conformal threshold with a single online Robbins–Monro (RM) offset to steer coverage toward a target level in real time. We benchmark TCP against GARCH, Historical Simulation, and a rolling Quantile Regression (QR) baseline across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold). Three results are consistent across assets. First, rolling QR yields the sharpest intervals but is materially under-calibrated (e.g., S&P 500: 83.2% vs. 95% target). Second, TCP (and TCP–RM) achieves near-nominal coverage across assets, with intervals that are wider than Historical Simulation in this evaluation (e.g., S&P 500: 5.21 vs. 5.06). Third, the RM update changes calibration and width only marginally at our default hyperparameters. Crisis‐window visualizations around March 2020 show TCP/TCP–RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage). A sensitivity study confirms robustness to window size and step‐size choices. Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
@@ 1.  INTRODUCTION
Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz, 1952; J.P. Morgan/Reuters, 1996). Yet, when markets enter turbulent regimes, traditional risk models often fall short. Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion, 2007; Dowd, 1998), assumptions that can systematically understate tail risk. Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar and Uryasev, 2000). Events like the 2008 global financial crisis (Brunnermeier, 2009) and the March 2020 stock-market crash (Mazur et al., 2021) exposed the fragility of parametric assumptions, especially in the tails.
The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques. Real-world returns are heteroskedastic, exhibit regime shifts, and often have heavy-tailed distributions (Cont, 2001). GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty. Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005). However, these guarantees rest on exchangeability, a condition often violated by time series with temporal dependencies. Several works extend conformal methods to sequential or temporally dependent settings (Xu and Xie, 2021; Stankeviciute et al., 2021; Gibbs and Candès, 2021), but a robust and practical solution for financial markets remains an open challenge.
@@
We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, and standard Quantile Regression (QR). Our results highlight a critical flaw in static ML approaches: while QR produces sharp intervals, it is poorly calibrated and systematically misses the 95% target. By contrast, TCP demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility to better align with the desired coverage rate. This adaptive property, which we visualize during the March 2020 crash, makes TCP a more reliable and principled framework for real-world risk management. A comprehensive sensitivity analysis further underscores the robustness of our framework to key hyperparameters.
@@ 2.  RELATED WORK
Traditional Financial Risk Models. Quantitative risk management traces its roots to portfolio theory (Markowitz, 1952) and was formalized around the VaR concept with J.P. Morgan’s RiskMetrics (J.P. Morgan/Reuters, 1996). In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion, 2007; Dowd, 1998; Glasserman, 2004). Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar and Uryasev, 2000). However, many traditional approaches underperform in crisis backtests (Kupiec, 1995; Christoffersen, 1998), highlighting a blind spot when tail events dominate.
Advanced Time-Series and Machine-Learning Models. To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle, 1982; Bollerslev, 1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson, 1991; Glosten et al., 1993). For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (III, 1975; McNeil and Frey, 2000), and VaR under heavy tails raises specific concerns (Danielsson and de Vries, 2000). On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker and Bassett, 1978), with ensemble and neural variants such as quantile regression forests (Meinshausen, 2006) and quantile networks (Taylor, 2000); deep LSTMs capture complex temporal dependencies (Fischer and Krauss, 2018). We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke et al., 2017).
Conformal Prediction Theory. Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk et al., 2005). Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs and Candès, 2021; Xu and Xie, 2021; Stankeviciute et al., 2021). Our work builds on these foundations with a practical adaptive conformal framework tailored to financial risk forecasting, combining local split-conformal calibration with an online Robbins–Monro offset for temporal adaptation.
@@ 3.  NOTATION
Variants used (terminology). We consider two closely related procedures. TCP denotes split–conformal prediction on a rolling window: the window $\mathcal{W}_{t}$ of size $w$ is split into a training slice $\mathcal{T}_{t}$ (size $w_{\mathrm{tr}}$) and a calibration slice $\mathcal{C}_{t}$ (size $w_{\mathrm{cal}}$), $w_{\mathrm{tr}}+w_{\mathrm{cal}}=w$. Quantile forecasters $\widehat{q}_{\alpha/2},\widehat{q}_{1-\alpha/2}$ are fit on $\mathcal{T}_{t}$ only and the two-sided split–conformal threshold $C_{t}$ is computed from $\mathcal{C}_{t}$, yielding $\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$. TCP-RM augments TCP with an online Robbins–Monro (RM) offset $C_{t}^{\mathrm{RM}}$ updated by $C_{t+1}^{\mathrm{RM}}=C_{t}^{\mathrm{RM}}+\gamma_{t}\{\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha\}$, with $\gamma_{t}=\gamma_{0}(1+\lambda t)^{-\beta}$. The effective threshold is $C_{t}^{\mathrm{eff}}=C_{t}+C_{t}^{\mathrm{RM}}$ and the interval becomes $\widehat{q}_{\alpha/2}(X_{t+1})-C_{t}^{\mathrm{eff}},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}^{\mathrm{eff}}]$. For comparability, the QR baseline is evaluated in the same walk-forward fashion using only the training slice ($|\mathcal{T}_{t}|=w_{\mathrm{tr}}$) and no calibration slice (rolling, out-of-sample).
@@ 4.  MATHEMATICAL BACKGROUND
This section lays the theoretical groundwork for our proposed method. We begin by formally defining the problem of prediction interval forecasting, then review the principles of classical conformal prediction and its limitations, and finally develop the core mathematical components of our Temporal Conformal Prediction (TCP) framework and its adaptive calibration mechanism.
@@ 5.1 Data and Feature Construction
  • Rolling Volatility: $\sigma_{t}=\sqrt{\frac{1}{20-1}\sum_{k=1}^{20}\big(r_{t-k}-\bar{r}_{t,20}\big)^{2}}\quad\text{(rolling sample std, ddof=1)}$ where $\bar{r}_{t,20}=\frac{1}{20}\sum_{k=1}^{20}r_{t-k}$.
@@
In implementation we use the rolling sample standard deviation via pandas (.rolling(20).std, ddof=1=1).
@@ 5.2.2 Benchmark Models
    $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\,\sigma_{t-1}^{2},\qquad(\omega,\alpha,\beta)=(10^{-6},\,0.05,\,0.9),$
@@ 5.3 Evaluation Framework
To provide a rigorous comparison, we must define clear success criteria. This subsection outlines the key performance metrics, empirical coverage and interval sharpness, that we use to evaluate and contrast the models. We use two primary metrics:
@@
Our analysis focuses on the trade-off between these two metrics, as a model is only useful if it is both well-calibrated (correct coverage) and sharp (informative).
@@ 6.  RESULTS AND DISCUSSION
We now assess Temporal Conformal Prediction (TCP) and our main contribution TCP-RM (TCP with a Robbins–Monro online calibration layer) against strong baselines (QR, GARCH, Historical Simulation) on three asset classes: S&P 500, BTC-USD, and Gold. The goal is to test whether the online calibration delivers near-nominal coverage under distribution shift while preserving sharpness.
Evaluation centers on two criteria: (i) empirical coverage versus the nominal target $1-\alpha=0.95$, and (ii) average interval width as a measure of sharpness. Table 1 summarizes full-sample performance; Figure 1 visualizes behavior during the COVID-19 crash. We also report a hyperparameter sensitivity study (window size $w$, stepsize $\gamma_{0}$) to assess robustness of TCP-RM.
@@
Figure 1: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for S&P 500 daily returns during the COVID-19 market crash (Feb–Apr 2020). Shaded bands show the interval; the red dots mark days where the realized return falls outside the 95% interval (miscoverage). TCP/TCP-RM bands widen rapidly into the March spike and contract in April, illustrating adaptiveness. Analogous crisis-window panels for BTC-USD and Gold appear in Appendix B.
@@
Three findings emerge clearly from Table 1 and the crisis-window visualization (Fig. 1):
(i) Calibration vs. sharpness trade-off. The rolling ML baseline (QR) yields the narrowest intervals across all assets but is materially under-calibrated relative to the 95% target: SP500 83.2%, BTC 85.5%, Gold 83.0%. This under-coverage is the price for sharpness (e.g., SP500 average width 3.311 versus TCP’s 5.210).
@@
(ii) TCP/TCP–RM achieve near-nominal coverage with wider bands than Historical Simulation. TCP coverage is close to the target on all assets (SP500 95.2%, BTC 95.4%, Gold 94.3%). Relative to Historical Simulation, the intervals are wider in this setting (SP500: 5.210 vs. 5.058; BTC: 20.888 vs. 18.058; Gold: 4.737 vs. 4.024), reflecting stricter calibration under elevated volatility.
@@
(iii) Robbins–Monro adds negligible change at these hyperparameters. TCP–RM’s coverage and width are essentially identical to TCP across assets (coverage deltas $\leq$ 0.0007; width deltas $\leq$ 0.002), indicating the online offset has minimal incremental effect here.
@@
GARCH under-covers consistently ($\approx$0.83–0.85) while being sharper than the calibrated methods, consistent with its conditional-Gaussian specification being slow to adapt to asymmetric tails during stress.
Crisis-window behavior. Figure 1 (SP500, Feb–Apr 2020) shows how each method behaves during the shock. The red dots mark days where the realized return falls outside the reported 95% prediction interval (miscoverage). TCP/TCP–RM bands widen abruptly into the March spike and contract as conditions stabilize, whereas QR remains comparatively tight and under-covers around extremes. GARCH adapts smoothly via its volatility recursion, and Historical Simulation reacts with a lag due to its long lookback. Overall, all methods temporarily under-cover at the break, but TCP/TCP–RM inflate bands right when volatility jumps and recover calibration thereafter. Analogous crisis-window panels for BTC-USD and Gold are reported in Appendix B.

### v5

@@ Abstract
We propose Temporal Conformal Prediction (TCP), a distribution-free framework for constructing well-calibrated prediction intervals in nonstationary time series. TCP couples a modern quantile forecaster with a split-conformal calibration layer on a rolling window and, in its TCP-RM variant, augments the conformal threshold with a single online Robbins-Monro (RM) offset to steer coverage toward a target level in real time. We benchmark TCP against GARCH, Historical Simulation, a rolling tree-based Quantile Regression (QR) model, a classical linear quantile regression baseline (QR-Linear), and an adaptive conformal method (ACI) across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold). Three results are consistent across assets. First, both QR and QR-Linear yield the sharpest intervals but are materially under-calibrated, and even ACI remains below the nominal 95% target in our full-sample backtests. Second, TCP (and TCP-RM) achieves near-nominal coverage across assets, with intervals that are wider than Historical Simulation in this evaluation (e.g., S&P 500: 5.21 vs. 5.06). Third, the RM update changes calibration and width only marginally at our default hyperparameters. Crisis-window visualizations around March 2020 show TCP/TCP-RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage). A sensitivity study confirms robustness to window size and step-size choices. Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
@@ 1.  INTRODUCTION
Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz1952; Morgan1996). Yet, when markets enter turbulent regimes, traditional risk models often fall short. Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion2007; Dowd1998), assumptions that can systematically understate tail risk. Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar2000). Events like the 2008 global financial crisis (Brunnermeier2009) and the March 2020 stock-market crash (Mazur2021) exposed the fragility of parametric assumptions, especially in the tails.
The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques. Real-world returns are heteroskedastic, exhibit regime shifts, and often exhibit heavy tails (Cont2001). GARCH-type models (Engle1982; Bollerslev1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty. Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk2005). However, these guarantees rest on exchangeability, a condition often violated by time series with temporal dependencies. Several works extend conformal methods to sequential or temporally dependent settings (Xu2021; Stankeviciute2021; Gibbs2021), but a robust and practical solution for financial markets remains an open challenge.
@@
We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, a tree-based Quantile Regression (QR) forecaster, a classical linear quantile regression baseline (QR-Linear), and an adaptive conformal method (ACI). Our results highlight a critical flaw in static ML approaches: both QR and QR-Linear produce sharp intervals but are systematically under-calibrated, and ACI improves calibration only partially. By contrast, TCP demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility so that empirical coverage stays close to the 95% target. This adaptive property, which we visualize during the March 2020 crash, makes TCP a more reliable and principled framework for real-world risk management. A comprehensive sensitivity analysis further underscores the robustness of our framework to key hyperparameters.
@@ 2.  RELATED WORK
Traditional Financial Risk Models. Quantitative risk management traces its roots to portfolio theory (Markowitz1952) and was formalized around the VaR concept with J.P. Morgan’s RiskMetrics (Morgan1996). In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion2007; Dowd1998; Glasserman2004). Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar2000). However, many traditional approaches underperform in crisis backtests (Kupiec1995; Christoffersen1998), highlighting a blind spot when tail events dominate.
Advanced Time-Series and Machine-Learning Models. To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle1982; Bollerslev1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson1991; Glosten1993). For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (Pickands1975; McNeil2000), and VaR under heavy tails raises specific concerns (Danielsson2000). On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker1978), with ensemble and neural variants such as quantile regression forests (Meinshausen2006) and quantile networks (Taylor2000); deep LSTMs capture complex temporal dependencies (Fischer2018). We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke2017).
Conformal Prediction Theory. Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk2005). Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs2021; Xu2021; Stankeviciute2021). Our work builds on these foundations with a practical adaptive conformal framework tailored to financial risk forecasting, combining local split-conformal calibration with an online Robbins–Monro offset for temporal adaptation.
Adaptive and Time-Series Conformal Prediction.
Recent research has addressed non-stationary environments by adapting the conformal guarantee. Gibbs2021 and zaffran2022adaptive introduce Adaptive Conformal Inference (ACI), which updates the target miscoverage rate ($\alpha_{t}$) online based on past coverage errors. While effective, we argue that in heavy-tailed financial settings, oscillating the target probability can lead to unstable interval widths; TCP-RM instead adapts the conformal threshold ($C_{t}$) directly via a Robbins-Monro update, offering a stable, scalar lever for width control. Other approaches, such as CPTD (lin2022temporal), leverage cross-sectional exchangeability across a panel of time series to restore validity. In contrast, TCP is designed for the univariate setting (single asset) where cross-sectional pooling (e.g., mixing Bitcoin and Gold) is inappropriate due to distinct, idiosyncratic asset dynamics.
@@ 3.  NOTATION
Variants used (terminology). We consider two closely related procedures. TCP applies split–conformal prediction on a rolling window, and TCP-RM augments TCP with a Robbins–Monro offset applied to the conformal threshold based on past coverage errors. Formal definitions and update equations are given in Section 4.4.
@@ 4.  MATHEMATICAL BACKGROUND
This section lays the theoretical groundwork for our proposed method. Our calibration mechanism is related in spirit to adaptive conformal procedures for nonstationary settings (zaffran2022adaptive; lin2022temporal) but differs in using a Robbins–Monro update on the conformal threshold while holding the nominal level fixed. We begin by formally defining the problem of prediction interval forecasting, then review the principles of classical conformal prediction and its limitations, and finally develop the core mathematical components of our Temporal Conformal Prediction (TCP) framework and its adaptive calibration mechanism.
@@ 4.3 Temporal Conformal Prediction (TCP)
Clarifications on Assumptions (A1–A4).
To make the role of the assumptions explicit in the analysis and in Theorem 4.2, we record the following clarifications:
  1. 1.
    Stochastic approximation framework. The Robbins-Monro update for the offset is analyzed using classical stochastic-approximation arguments under mixing and bounded-variance noise; our assumptions instantiate those conditions in this setting ((Kushner2003)).
  2. 2.
    TCP-RM guarantees. Under (A1)–(A4), the time-average coverage converges to the target level and the effective threshold converges to a fixed point (Theorem 4.2).
  3. 3.
    TCP baseline (split-conformal). TCP uses split-conformal calibration within a rolling window; this is a local working approximation to exchangeability, and we do not claim global finite-sample validity for the full time series.
  4. 4.
    Assumption verification in practice. We empirically probe the mixing/independence assumptions using standard risk diagnostics (Kupiec1995 and Christoffersen1998 backtests); the lack of serial dependence in exceedances supports the validity of these conditions on our financial datasets.
@@ 5.1 Data and Feature Construction
  • Rolling Volatility:
    $\sigma_{t}$ $=\sqrt{\frac{1}{20-1}\sum_{k=1}^{20}(r_{t-k}-\bar{r}_{t,20})^{2}},$ (3)
    $\bar{r}_{t,20}$ $=\frac{1}{20}\sum_{k=1}^{20}r_{t-k}.$ (4)
    Here $\sigma_{t}$ is the rolling sample standard deviation of the past 20 returns, computed with denominator 20120-1 (in pandas this corresponds to .rolling(20).std(ddof=1)).
@@ 5.2.2 Benchmark Models
    Linear Quantile Regression (QR-Linear): Classical linear quantile regression trained with the pinball loss at $\alpha/2$ and $1-\alpha/2$, using the same features and the same rolling, out-of-sample protocol as QR (window $w_{\mathrm{tr}}{=}192$; no calibration slice). This baseline is included in the full-sample numerical benchmarks (Table 1), but we omit it from the crisis-window visualizations to avoid clutter; its behavior is intermediate between QR and ACI.
  3. 3.
    Adaptive Conformal Inference (ACI): A rolling implementation of Gibbs2021 that uses the same window split as TCP ($w_{\mathrm{tr}}=192$ training, $w_{\mathrm{cal}}=60$ calibration). However, instead of updating the conformal threshold $C_{t}$ (like TCP-RM), ACI updates the target miscoverage rate $\alpha_{t}$ online based on the most recent coverage error (step size $\eta=0.05$).
  4. 4.
@@
    $\sigma_{t}^{2}$ $=\omega+\alpha r_{t-1}^{2}+\beta\,\sigma_{t-1}^{2},$ (5)
    $(\omega,\alpha,\beta)$ $=(10^{-6},\,0.05,\,0.9),$ (6)
@@ 5.3 Evaluation Framework
To provide a rigorous comparison, we define clear success criteria for all models. Our primary metrics are (i) marginal empirical coverage of the prediction intervals and (ii) average interval width (sharpness). In addition, we apply standard VaR backtests on the 5% lower tail implied by the 95% two-sided intervals.
We use two core summary metrics:
@@
VaR-style backtests.
For each model and asset, we also perform standard 1-day-ahead VaR backtests on the 5% lower tail. Specifically, we report: (i) the unconditional coverage test of Kupiec1995 (LRuc), (ii) the independence test of Christoffersen1998 (LRind), and (iii) their sum $_{\mathrm{cc}}{=}LR$$_{\mathrm{uc}}{+}$LRind for conditional coverage. Each test yields a likelihood-ratio statistic and a $p$-value, together with the number of exceedances $x$ (days where the realized return falls below the nominal 5% VaR) out of $n$ out-of-sample predictions.
Our empirical analysis therefore evaluates models along two dimensions, i.e., global calibration–sharpness, via empirical coverage and average width over the full sample, and risk-management diagnostics in the left tail, via the Kupiec/Christoffersen backtests.
@@ 6.  RESULTS AND DISCUSSION
We now assess Temporal Conformal Prediction (TCP) and our main contribution TCP-RM (TCP with a Robbins–Monro online calibration layer) against a set of baselines (ACI, QR, QR-Linear, GARCH, Historical Simulation) on three asset classes: S&P 500, BTC-USD, and Gold. The goal is to test whether the online calibration delivers near-nominal coverage under distribution shift while preserving sharpness.
Evaluation centers on two criteria: (i) empirical coverage versus the nominal target $1-\alpha=0.95$, and (ii) average interval width as a measure of sharpness. Table 1 summarizes full-sample performance; Figure 1 visualizes behavior during the COVID-19 crash. We also report a hyperparameter sensitivity study (window size $w$, stepsize $\gamma_{0}$) to assess robustness of TCP-RM. QR-Linear appears in the full-sample benchmark table but is omitted from the crisis-window plots and window-level summary tables to keep the figures legible; numerically, its coverage and width lie between QR and ACI.
@@
Figure 1: A comparison of 95% prediction intervals from six models (TCP, TCP-RM, QR, GARCH, Hist, ACI) for S&P 500 daily returns during the COVID-19 market crash (Feb–Apr 2020). Shaded bands show the interval; the red dots mark days where the realized return falls outside the 95% interval (miscoverage). TCP/TCP-RM bands widen rapidly into the March spike and contract in April, while ACI maintains relatively narrow bands and under-covers more frequently. Analogous crisis-window panels for BTC-USD and Gold appear in Appendix E.
@@
Three findings emerge from Table 1 and the crisis-window visualization (Fig. 1):
(i) Calibration vs. sharpness trade-off across learning baselines. The rolling ML baselines (QR, QR-Linear, ACI) yield substantially narrower intervals than TCP/TCP-RM, but remain under-calibrated relative to the 95% target, which is consistent with the VaR backtests where only TCP and TCP-RM are not rejected. Among these baselines, QR-Linear and ACI partially improve calibration over tree-based QR at the price of somewhat wider bands, yet all three still fall short of nominal coverage.
@@
(ii) TCP/TCP–RM achieve near-nominal coverage with wider bands than Historical Simulation. TCP coverage is close to the 95% target on all assets (SP500 95.2%, BTC-USD 95.4%, Gold 94.3%). Relative to Historical Simulation, TCP/TCP-RM intervals are systematically wider (SP500: 5.21 vs. 5.06; BTC-USD: 20.89 vs. 18.06; Gold: 4.74 vs. 4.02), reflecting stricter calibration under elevated volatility. Historical Simulation attains high but slightly sub-nominal coverage (93–94%) with narrower bands, consistent with its reliance on a long empirical window that smooths abrupt regime changes.
@@
(iii) Robbins–Monro adds negligible change at these hyperparameters. TCP–RM’s coverage and width are essentially identical to TCP across assets (coverage deltas $\leq 0.0007$; width deltas $\leq 0.01$), indicating that the additional Robbins–Monro offset has minimal incremental effect at the chosen $(w,\gamma_{0})$ settings.
@@
GARCH under-covers consistently (about 82–85% coverage) while being sharper than the calibrated methods, consistent with its conditional-Gaussian specification being slow to adapt to asymmetric tails during stress.
Crisis-window coverage.
To stress-test the models, we also examine a 55-day “COVID crisis” window from February–April 2020 for each series. For this sub-period we compute the empirical coverage and average width of the 95% prediction intervals. Table 2 reports the results for the S&P 500; the corresponding tables for BTC-USD and Gold are given in Appendix B.
For the S&P 500, TCP and TCP–RM both achieve 80% coverage with very similar average widths (11.94 and 11.95, respectively). GARCH attains the same coverage (80%) with a narrower interval (average width 9.66), while the rolling quantile regression and historical benchmark produce even sharper intervals (average widths 9.13 and 5.91) at the cost of substantially lower coverage (65.45% each). ACI is the most aggressive, with the narrowest intervals (average width 4.63) but only 56.36% coverage. Thus, during the sharpest part of the COVID shock, all methods under-cover, but TCP and TCP-RM remain competitive with GARCH in terms of the calibration-sharpness trade-off, while clearly dominating ACI and the nonparametric baselines in coverage.
Table 2: Crisis-window empirical coverage and average width of 95% prediction intervals for the S&P 500 during the COVID shock (Feb–Apr 2020, $n=55$ days).
Model Coverage Avg. width $n$
TCP 0.8000 11.9371 55
TCP-RM 0.8000 11.9511 55
QR (rolling) 0.6545 9.1315 55
GARCH 0.8000 9.6571 55
Hist 0.6545 5.9058 55
ACI 0.5636 4.6307 55
TCP versus TCP–RM.
Figure 2 overlays TCP and TCP–RM intervals with the realized returns during the COVID window for the S&P 500; analogous plots for BTC-USD and Gold are provided in Appendix C. The bands for TCP and TCP–RM are almost indistinguishable in all three series, and this is confirmed quantitatively.
Over the COVID window, TCP and TCP–RM have identical empirical coverage for each asset (0.8000 for S&P 500, 0.8909 for BTC-USD, and 0.8727 for Gold), and their average interval widths differ only at the third decimal place: 11.94 vs 11.95 for S&P 500, 24.28 vs 24.27 for BTC-USD, and 6.81 vs 6.81 for Gold. Over the full sample, the two methods also remain nearly identical: for S&P 500, coverage is 0.9523 for TCP and 0.9530 for TCP–RM with average widths 5.2097 vs 5.2171 and interval scores 7.1168 vs 7.1157; for BTC-USD, both methods achieve coverage 0.9537 with widths 20.8879 vs 20.8859 and interval scores 27.9846 vs 27.9851; for Gold, both achieve coverage 0.9427 with widths 4.7368 vs 4.7438 and interval scores 6.3246 vs 6.3246. In other words, the Robbins–Monro adaptation layer leaves the marginal predictive performance of TCP essentially unchanged while preserving its good calibration properties.
Figure 2: S&P 500: TCP vs TCP–RM 95% prediction intervals during the COVID crisis window (Feb–Apr 2020).